eviant:

A model of a car is 6.3 in. long. The scale of the model to that actual car is 1:30. What is the length of the actual car to the nearest foot?

eviant:

@Vocaloid 16 ft?

jimthompson5910:

"The scale of the model to that actual car is 1:30" meaning that the actual car's length is 30 times greater than that of its scale model counter part real length = 30*(scale model length) real length = 30*(6.3 inches) real length = 189 inches 189 inches = (189 in)*(1 ft/12 in) = (189/12) ft = 15.75 ft which rounds to 16 feet You have the correct answer. Nice work.